OpenStudy (xxbatmanevoxx):

WILL FAN AND MEDAL A set of data with 100 observations has a mean of 267. There are six outliers in the data set, which have a mean of 688. If the six outliers are removed, what is the mean of the new data set?

OpenStudy (xxbatmanevoxx):

@Samara8954

OpenStudy (caiti2024):

X*94 +688x6 = 267x100 94x+4128=26700 94X=22572/94 X=240.128

OpenStudy (xxbatmanevoxx):

OMG thx

OpenStudy (caiti2024):

No problem. :)